[ The sum of the ages of the father and son at present is 45 years. If both live on until the son's age becomes equal to the father's present age,the sum of their ages will be 95 years,find their present ages?
Thank you

Answer by TimothyLamb(4379)

f + s = 45
f = s + X
f - s = X
(f + X) + (s + X) = 95
---
(f + X) + (s + X) = 95
f + s + 2X = 95
45 + 2X = 95
2X = 50
X = 25
---
f + s + 2X = 95
(s + X) + s + 50 = 95
(s + 25) + s + 50 = 95
2s + 75 = 95
2s = 20
---
s = 10
f = 35
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php